OpenStudy (anonymous):

Change the given equation into slope-intercept form. 2x + 5y = -10?

OpenStudy (anonymous):

\[2x+5y=-10\] \[-2x\] \[5y=-2x-10\] \[\div5\] \[y=\frac{ -2 }{ 5 }-2\]

OpenStudy (anonymous):

the fraction is -2/5x

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

any time

OpenStudy (anonymous):

can you explain though

OpenStudy (anonymous):

you take 2x and subtract it to the other side, then divide everything by 5, and you get the answer \[y=\frac{ -2 }{ 5}x-2\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

The inverse of the function y = 2x + 8 is y= x -8/ 2 True or False